OpenStudy (anonymous):

A farmer has 1,300 acres of land on which he grows corn, wheat, and soybeans. It costs $45 per acre to grow corn, $60 to grow wheat, and $50 to grow soybeans. Because of market demand the farmer will grow twice as many acres of wheat as of corn. He has allocated $69,050 for the cost of growing his crops. How many acres of each crop should he plant? corn ____ acres wheat ____ acres soybeans _____ acres

OpenStudy (gw2011):

I got the following answers: corn=270 acres wheat=540 acres soybeans=490 acres Here is how I arrived at these answers: Let c=corn Let w=wheat=2c Let s=soybeans 45c+120c+50s=69,050 (60 times 2=120c) c+2c+s=1300 165c+50s=69,050 3c+s=1300 Solving these two equations by eliminating s (multiply second equation by 50) 165c+50s=69,050 150c+50s=65,000 15c=4050 c=270 acres w=540 acres (270 times 2) s=490 acres (1300-270-540)